OpenStudy (anonymous):

If the area of a circle is 8pi, what is the length of it's diameter? Will fan/medal:)

jimthompson5910 (jim_thompson5910):

\[\Large A = \pi*r^2\] \[\Large 8\pi = \pi*r^2\] \[\Large r = ???\]

OpenStudy (anonymous):

I've tried that I just can't seem to figure out the first step I need to take!

jimthompson5910 (jim_thompson5910):

how do you get r^2 all by itself

OpenStudy (anonymous):

Somehow get rid of the pi, but I'm not sure how to?

jimthompson5910 (jim_thompson5910):

why not divide both sides by pi

OpenStudy (anonymous):

I will try that

jimthompson5910 (jim_thompson5910):

that will get rid of pi on both sides because pi/pi = 1

jimthompson5910 (jim_thompson5910):

after r^2 is isolated you have to isolate r itself

OpenStudy (anonymous):

so square root r^2?

jimthompson5910 (jim_thompson5910):

square root both sides

OpenStudy (anonymous):

I got 2.8284721=r, but the only answer choices I have are a. 4 b. 8 c. \[2\sqrt{2}\] d. \[4\sqrt{2}\]

OpenStudy (anonymous):

I don't know if I have to convert to a radical.. If so, I don't know how to.

jimthompson5910 (jim_thompson5910):

Simplify \[\Large \sqrt{8}\]

OpenStudy (anonymous):

\[2\sqrt{4}?\]

OpenStudy (anonymous):

vice versa

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{8} = \sqrt{4*2} = ???\]

OpenStudy (anonymous):

\[4\sqrt{2}\] i mixed it up on accident

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{8} = \sqrt{4*2} = \sqrt{4}*\sqrt{2} = ???\]

OpenStudy (anonymous):

\[2\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

That's the radius

jimthompson5910 (jim_thompson5910):

We want the diameter

jimthompson5910 (jim_thompson5910):

\[\Large r = 2\sqrt{2}\] \[\Large d = ???\]

OpenStudy (anonymous):

urggg. so its not \[4\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

it is, I'm getting there

jimthompson5910 (jim_thompson5910):

I want to make sure you know how to get there

OpenStudy (anonymous):

Okay, I've never been good at radicals.. Sorry!

jimthompson5910 (jim_thompson5910):

that's ok, just keep practicing and you'll get better

OpenStudy (anonymous):

Yeah I just need A LOT of practice :) But you explain things to me very well!

jimthompson5910 (jim_thompson5910):

Glad some of it is starting to make sense

OpenStudy (anonymous):

Me too, trust me.. :)